OpenStudy (anonymous):

Odell tried to solve the equation 6(x+7)=4x-3 . He recorded the following steps. 6(x+7)=4x-3 6x=4x-10 2x=-10 x=-5 When Odell checked his solution it did not work. What was his mistake? Explain. Show the correct solution.

OpenStudy (ybarrap):

6(x + 7) = 4x - 3 6x + 42 = 4x - 3 6x = 4x - 3 - 42 = 4x - 45 Ans. Odell made his mistake in the second step.

OpenStudy (turingtest):

this is already posted and solved above